Find the accumulated amount after 4 years if $1000 is invested at 6% per year compounded (a) daily (assume a 365-day year) and (b) continuously...Find the accumulated amount after 4 years if $1000 is invested at 6% per year compounded (a) annually, (b) semiannually, (c) quarterly, (d) monthly, and (e) daily.


1
Expert's answer
2021-06-02T13:57:00-0400

1.

(a) Compound interest


"A=P(1+\dfrac{r}{n})^{nt}"

Daily: "n=365"


"A=1000(1+\dfrac{0.06}{365})^{365(4)}=\$1271.22"

(b) Continuously Compounded Interest


"A=Pe^{rt}=1000e^{0.06(4)}=\$1271.25"

2.

Compound interest


"A=P(1+\dfrac{r}{n})^{nt}"

(a) Annually "n=1"


"A=1000(1+\dfrac{0.06}{1})^{1(4)}=\$1262.48"

(b) Semiannually "n=2"


"A=1000(1+\dfrac{0.06}{2})^{2(4)}=\$1266.77"


(c) Quarterly "n=4"


"A=1000(1+\dfrac{0.06}{4})^{4(4)}=\$1268.99"



(d) Monthly "n=12"


"A=1000(1+\dfrac{0.06}{12})^{12(4)}=\$1270.49"



(e) Daily: "n=365"


"A=1000(1+\dfrac{0.06}{365})^{365(4)}=\$1271.22"